# Adaptive Off-Policy Inference for M-Estimators Under Model Misspecification

### v1

@@ 2 Problem Setup
$s_{t,\theta}=\sum_{a=1}^{K}\pi_{e}(A_{t}=a|X_{t})\left(\dot{m}_{\theta}(a,X_{t},f_{t}(a,X_{t}))+\mathbbm{1}_{A_{t}=a}\frac{\left(\dot{m}_{\theta}(a,X_{t},Y_{t})-\dot{m}_{\theta}(a,X_{t},f_{t}(A_{t},X_{t})\right)}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}\right).$ (6)
In order to limit the variance of this quantity, we will instead consider a second estimator $\hat{\theta}_{T}$ defined as the solution to $\sum_{t=1}^{T}\Sigma_{t}^{-1}s_{t,\theta}=0$ for specially constructed matrices $\Sigma_{t}\in\mathbb{R}^{d}$ designed to normalize the variance of $s_{t,\theta}$. However, the construction of $\Sigma_{t}$ will require consistent estimates of $\theta^{\star}$ so our procedure will involve a two-step process where $\tilde{\theta}_{T}$ is used to construct the stabilizing matrices $\Sigma_{t}$ and then a central limit theorem is proved for $\hat{\theta}_{T}$.
@@ 3 Main Results
Our approach will be to construct two sets of weights based on different filtrations. As before, $w_{t}\in\sigma(X_{t},\mathcal{H}_{t-1})$ is restricted to being the inverse propensity weights $\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}(A_{t}|X_{t},\mathcal{H}_{t-1})}$ to ensure that $\hat{\theta}_{T}$ is unbiased for $\theta^{\star}$. Simultaneously, we consider $\Sigma_{t}\in\sigma(\mathcal{H}_{t-1})$ that will allow us to adjust the second moment, while preserving the first moment of the score function at $\theta^{\star}$ to be 0. By assumption, $\mathbb{E}\left[s_{t,\theta^{\star}}\right]=0$. Define a martingale difference sequence (MDS) $Z_{t}:=\Sigma_{t}^{-1/2}s_{t,\theta^{\star}}$. Then,
@@
the fact that $\Sigma_{t}^{-1/2}\in\sigma(\mathcal{H}_{t-1})$. A natural choice is to choose $\Sigma_{t}:=V_{t,\theta^{\star}}=\mathbb{E}_{\mathcal{P},p_{t}}\left[s_{t,\theta^{\star}}s_{t,\theta^{\star}}^{T}|\mathcal{H}_{t-1}\right]$.
By re-weighting the score function to stabilize the variance but requiring that the weights be determined from a coarser filtration than what was used for constructing $w_{t}$, we are able to preserve the first-order conditions from 3. Of course, $V_{t,\theta^{\star}}$ is unknown in practice and will need to be replaced with an estimate $\hat{V}_{t}$. We proceed with the remainder of this section assuming the existence of such an estimate but note that it is a non-trivial task as we only have a single observation corresponding to each $t$. The key idea for constructing an estimator is to leverage the potential outcomes framework of 1, which we discuss in Section 4.
@@
The first three derivatives of $m_{\theta}$(x,a,y) exist for every $\theta\in\Theta$, every $a\in\mathcal{A}$ and every $(x,y)$ in the joint support of $\mathcal{P}$.
@@
Let $\theta^{\star}$ be defined as in Equation 2 and $\Theta$ be defined as in 2. We further assume that at every $t\in[T]$,
@@
    For all sufficiently large $T$, for any $\epsilon>0$, there exists constants $\delta_{1},\delta_{2}>0$ such that
@@
The first two conditions of this assumption ensures that the quantity of interest $\theta^{\star}$ is the solution to a maximization problem and can be found by evaluating the critical points of $m_{\theta}$. The third condition mirrors classical assumptions to demonstrate the consistency of $Z$ estimate and $M$ estimators (van der Vaart, 2000). Having a well-separated solution ensures the uniqueness of $\theta^{\star}$ and allows estimators constructed from finite samples to converge appropriately. Note that we assume that $\theta^{\star}$ is well separated both as a maximizer of $m_{\theta}$ and as a root of the score equation because we consider estimators defined in both of these ways within our procedure. A special case of the third condition is when $m_{\theta}$ is a continuously differentiable concave function with the maximum obtained at $\theta^{\star}$ .
@@
Define the set of functions $\mathcal{M}_{\Theta}=\{m_{\theta}(X_{t},A_{t},Y_{t}):\theta\in\Theta\}$ and $\dot{\mathcal{M}}_{\Theta}=\{c^{T}\dot{m}_{\theta}(X_{t},A_{t},Y_{t}):\theta\in\Theta,\left\lVert c\right\rVert<1\}$. We assume that for all $\epsilon>0$, the bracketing numbers $N_{[]}(\epsilon,\mathcal{M}_{\Theta},L_{2}(\mathcal{P},\pi_{e}))<\inftyand$ $N_{[]}(\epsilon,\dot{\mathcal{M}}_{\Theta},L_{2}(\mathcal{P},\pi_{e}))<\infty$.
5 limits the complexity of the function class $\{m_{\theta}:\theta\in\Theta\}$ so that a martingale law of large numbers can be applied which is required when constructing an argument for the consistency of $\hat{\theta}_{T}$ and $\tilde{\theta}_{T}$. We note that Zhang et al. (2021) provides an intuitive sufficient condition to check to ensure that 5 is true. They show that whenever $m_{\theta}$ is Lipschitz in the sense that,
@@
for some function $h$ such that $\mathbb{E}_{\mathcal{P},p_{e}}[h(X_{t},A_{t},Y_{t})^{2}]<m$ for some constant $m$, then $\mathcal{M}_{\theta}$ will have finite btracketing numbners for all $\epsilon>0$. We place a similar assumption on the complexity of the model classes.
@@
The predictive model $f_{t}\in\sigma(\mathcal{H}_{t-1})$ and is restricted to belong to a function class $\mathcal{F}$. Furthermore, we assume that there exists a function $u_{\mathcal{F}}$ such that $\mathbb{E}[u_{\mathcal{F}}(X_{t},a,f_{t}(a,X_{t}))^{2}]<m_{u}$ for all $a\in\mathcal{A}$ and some constant $m_{u}\in\mathbb{R}$ such that
@@
for all $a\in\mathcal{A}$, $x$ in the support of $\mathcal{P}$ and $f_{t}\in\mathcal{F}$.
Note that a primary difference between 5 and 6 is that 6 is written in terms of expectations relative to the actual policy $\pi_{t}$ rather than $\pi_{e}$. This is necessary because 6 bounds the variance of a term related to $f_{t}$, which is dependent on history whereas 5 is bounding the complexity of a quantity that is independent of history. These assumptions allow us to prove a key lemma needed to ensure the consistency of $\hat{\theta}_{T}$ and $\tilde{\theta}_{T}$,.
@@
Then under Assumptions 1-4 and 6,
$\sup_{\theta\in\Theta}\frac{1}{T}\sum_{t=1}^{T}R_{t}(\theta)-\mathbb{E}[R_{t}(\theta)|\mathcal{H}_{t-1}]\overset{p}{\to}0.$
@@
and $\mathbb{E}_{\mathcal{P},p_{e}}\left[u_{m}(X_{t},A_{t},Y_{t})^{2}\right]$ is bounded.
@@
Let $\theta^{\star}$ be defined as Equation 2 and $s_{t,\theta}$ defined as in Equation 6. Assume $V_{t,\theta^{\star}}:=\mathbb{E}_{\mathcal{P},p_{t}}\left[s_{t,\theta^{\star}}s_{t,\theta^{\star}}^{T}|\mathcal{H}_{t-1}\right]$ is almost surely invertible and has eigenvalues bounded above and below by constants $\delta_{\text{min}},\delta_{\text{max}}$. Assume that there exists a sequence of invertible random matrices $\{\hat{V}_{t}\}_{t=1}^{T}$ adapted to the filtration $\sigma(\mathcal{H}_{t-1})$ such that $\left\lVert\hat{V}_{t}^{-1/2}-V_{t,\theta^{\star}}^{-1/2}\right\rVert_{\text{op}}\overset{p}{\to}0$. Let $\hat{\theta}_{T}$ be any estimator such that $\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\hat{\theta}_{T}}=o_{p}(1)$. Then under Assumptions 1-6,
@@
Note that etting $\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\hat{\theta}_{T}}=o_{p}(1)$ rather than exactly zero ensures that the results are still valid when using an approximate root finding algorithm instead of an exact root. As an illustrative example, we first demonstrate the application of Theorem 1 in the case of a linear model.
Example 2 (Linear Model).
@@
Consider the (correctly specified) linear model $Y_{t}=\theta^{T}X_{t}+\epsilon_{t}$, $\epsilon_{t}\sim N(0,\sigma^{2})$, where $X_{t}\in\sigma(\mathcal{H}_{t-1})$ is dependent on user choices based on a shared history $\mathcal{H}_{t-1}:=\{X_{1},Y_{1},...,X_{t-1},Y_{t-1}\}$. Letting $\Sigma_{T}:=\sum_{t=1}^{T}X_{t}X_{t}^{T}$, Lai and Wei (1982) show that a sufficient condition for the OLS estimator $\tilde{\theta}_{T}:=\Sigma_{T}^{-1}\sum_{t=1}^{T}X_{t}Y_{t}$ to converge to $\theta$ almost surely is that $\lambda_{\text{min}}(\Sigma_{T})\overset{\text{a.s.}}{\to}0$ and $\frac{\log\lambda_{\text{max}}(\Sigma_{T})}{\lambda_{\text{min}}(\Sigma_{T})}\overset{\text{a.s.}}{\to}0$. The requirement for asymptotic normality, however, is much stronger. They require that a deterministic series of matrices $\{B_{T}\}_{T=1}^{\infty}$ exist such that $B_{T}^{-1}(\sum_{t=1}^{T}X_{t}X_{t}^{T})^{1/2}\overset{p}{\to}I_{p}$.
@@
Suppose that we apply Theorem 1 in the case where $\tilde{\theta}_{T}$ is consistent but the asymptotic normality condition does not apply. We can still form a confidence interval as follows. For simplicity, we assume that no predictive models are trained so the score function corresponds to the ordinary least squares loss. In this case, $V_{t}=\sigma_{t}^{-1}\mathrm{Var}\left(X_{T}^{T}(Y-\theta^{T}X_{t})|\mathcal{H}_{t-1}\right)=\sigma_{t}^{2}\mathbb{E}\left[X_{t}X_{t}^{T}|\mathcal{H}_{t-1}\right]$ for $\sigma_{t}^{2}:=\text{Var}\left(Y_{t}|X_{t}\right)$ and $\dot{s}_{t,\theta}=X_{t}X_{t}^{T}$. We then have that,
@@
$\hat{\theta}_{T}=\left(\sigma_{t}^{-1}\sum_{t=1}^{T}\mathbb{E}\left[X_{t}X_{t}^{T}|\mathcal{H}_{t-1}\right]^{-1/2}X_{t}X_{t}^{T}\right)^{-1}\left(\sum_{t=1}^{T}\sigma_{t}^{-1}\mathbb{E}\left[X_{t}X_{t}^{T}|\mathcal{H}_{t-1}\right]^{-1/2}X_{t}Y_{t}\right).$
Forming confidence intervals using Theorem 1 yields,
@@
$\sum_{t=1}^{T}\sigma_{t}^{-1}\mathbb{E}\left[X_{t}X_{t}^{T}|\mathcal{H}_{t-1}\right]^{-1/2}X_{t}X_{t}^{T}\left(\hat{\theta}_{T}-\theta\right)\overset{d}{\to}N(0,I_{p}).$
In the case where $\mathbb{E}\left[X_{t}X_{t}^{T}|\mathcal{H}_{t-1}\right]$ is independent of $\mathcal{H}_{t-1}$, then $\frac{1}{T}\sum_{t=1}^{T}X_{t}X_{t}^{T}$ becomes a consistent estimate for $\mathbb{E}\left[X_{t}X_{t}^{T}\right]$. When $\sigma_{t}$ is constant across $t$, this reduces to the ordinary least squares solution (i.e. $\tilde{\theta}_{T}=\hat{\theta}_{T}$) and the confidence intervals become standard. When $\sigma_{t}$ is not constant, $\hat{\theta}$ corresponds to a weighted least squares estimator. When $X_{t}\in[0,1]^{d}$ consists of $d$ indicator variables, then $\mathbb{E}\left[X_{t}X_{t}^{T}|\mathcal{H}_{t-1}\right]=\sigma_{t}^{2}\text{diag}\left(\mathbb{P}\left(X_{t,j}=1|\mathcal{H}_{t-1}\right)^{-1}\right),$ which aligns exactly with the inverse propensity weighted solution for (noncontextual) bandits.
Of course, the key challenge in applying Theorem 1 is to find a reliable method of estimating $\{\hat{V}_{t}\}_{t=1}^{T}$ which we tackle in the next section.

### v2

@@ 2 Problem Setup
$s_{t,\theta}=\sum_{a=1}^{K}\pi_{e}(A_{t}=a|X_{t})\left(\dot{m}_{\theta}(X_{t},a,f_{t}(X_{t},a))+\mathbbm{1}_{A_{t}=a}\frac{\left(\dot{m}_{\theta}(X_{t},a,Y_{t})-\dot{m}_{\theta}(X_{t},a,f_{t}(X_{t},a)\right)}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}\right).$ (7)
In order to limit the variance of this quantity, we will instead consider a second estimator $\hat{\theta}_{T}$ defined as the (approximate) root of $\sum_{t=1}^{T}\Sigma_{t}^{-1}s_{t,\theta}$ for specially constructed matrices $\Sigma_{t}\in\mathbb{R}^{d}$ designed to normalize the variance of $s_{t,\theta}$. However, the construction of $\Sigma_{t}$ will require consistent estimates of $\theta^{\star}$ so our procedure will involve a two-step process where $\tilde{\theta}_{T}$ is used to construct the stabilizing matrices $\Sigma_{t}$ and then a central limit theorem is proved for $\hat{\theta}_{T}$.
@@ 3 Main Results
Our approach will be to construct two sets of weights based on different filtrations. As before, $w_{t}\in\sigma(X_{t},\mathcal{H}_{t-1})$ is restricted to being the inverse propensity weights $\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}(A_{t}|X_{t},\mathcal{H}_{t-1})}$ to ensure that $\hat{\theta}_{T}$ is unbiased for $\theta^{\star}$. Similarly, we consider $\Sigma_{t}\in\sigma(\mathcal{H}_{t-1})$ that will allow us to adjust the second moment of $s_{t,\theta}$, while preserving the first moment of the score function in $\theta^{\star}$ to be 0. By assumption, $\mathbb{E}\left[s_{t,\theta^{\star}}|\mathcal{H}_{t-1}\right]=0$. Define a martingale difference sequence (MDS) $Z_{t}:=\Sigma_{t}^{-1/2}s_{t,\theta^{\star}}$. Then,
@@
by the fact that $\Sigma_{t}^{-1/2}\in\sigma(\mathcal{H}_{t-1})$. A natural choice is to choose $\Sigma_{t}$ as
$V_{t,\theta^{\star}}=\mathbb{E}_{\mathcal{P},\pi_{t}}\left[s_{t,\theta^{\star}}s_{t,\theta^{\star}}^{T}|\mathcal{H}_{t-1}\right].$
By re-weighting the score function to stabilize the variance but requiring that the weights be determined from a coarser filtration than what was used for constructing $w_{t}$, we are able to preserve the first-order conditions that will be stated in 3. Of course, $V_{t,\theta^{\star}}$ is unknown in practice and will need to be replaced with an estimate $\hat{V}_{t}$. We proceed with the remainder of this section assuming the existence of such an estimate but note that it is a non-trivial task as we only have a single observation corresponding to each $t$. The key idea for constructing an estimator is to leverage the potential outcomes framework of 1, which we discuss in Section 4.
@@
The first three derivatives of $m_{\theta}(x,a,y)$ with respect to $\theta$ exist for every $\theta\in\Theta$, every $a\in\mathcal{A}$ and every $(x,y)$ in the joint support of $\mathcal{P}$.
@@
Let $\theta^{\star}$ be defined as in Equation 3 and $\Theta\subset\mathbb{R}^{d}$ be a bounded parameter space such that 2 holds. We further assume that at every $t\in[T]$,
@@
    For any $\epsilon>0$, there exists constants $\delta_{1},\delta_{2}>0$ such that
@@
  4. 4.
    The first four moments of $m_{\theta}(X_{t},A_{t},Y_{t})$, $\dot{m}_{\theta}(X_{t},A_{t},Y_{t})$, and $\ddot{m}_{\theta}(X_{t},A_{t},Y_{t})$ are bounded with respect to $\mathcal{P},\pi_{e}$.
The first two conditions of this assumption ensure that the quantity of interest $\theta^{\star}$ is the solution to a maximization problem and can be found by evaluating the critical points of $m_{\theta}$. The third condition mirrors classical assumptions to demonstrate the consistency of $Z$ estimators and $M$ estimators (van der Vaart, 2000). Having a well-separated solution ensures the uniqueness of $\theta^{\star}$ and allows estimators constructed from finite samples to converge appropriately. Note that we assume that $\theta^{\star}$ is well separated both as a maximizer of $m_{\theta}$ and as a root of the score equation because we consider estimators defined in both of these ways within our procedure. A special case of the third condition is when $m_{\theta}$ is a continuously differentiable concave function with the maximum obtained at $\theta^{\star}$. The fourth condition is used to ensure that Lindeberg conditions are achieved when invoking the martingale central limit theorem and is analogous to conditions in classical proofs of the normality of $M$-estimators (van der Vaart, 2000).
@@
Define the set of functions $\mathcal{M}_{\Theta}=\{m_{\theta}(X_{t},A_{t},Y_{t}):\theta\in\Theta\}$ and $\dot{\mathcal{M}}_{\Theta}=\{c^{T}\dot{m}_{\theta}(X_{t},A_{t},Y_{t}):\theta\in\Theta,\left\lVert c\right\rVert\leq 1\}$. We assume that for all $\epsilon>0$, the bracketing numbers $N_{[]}(\epsilon,\mathcal{M}_{\Theta},L_{2}(\mathcal{P},\pi_{e}))<\infty$ and $N_{[]}(\epsilon,\dot{\mathcal{M}}_{\Theta},L_{2}(\mathcal{P},\pi_{e}))<\infty$.
5 limits the complexity of the function class $\{m_{\theta}:\theta\in\Theta\}$ so that a martingale law of large numbers can be applied which is required when constructing an argument for the consistency of $\hat{\theta}_{T}$ and $\tilde{\theta}_{T}$. We note that Zhang et al. (2021) provides an intuitive sufficient condition to check to ensure that 5 is true. They show that whenever $m_{\theta}$ is Lipschitz in the sense that
@@
for some function $h$ such that $\mathbb{E}_{\mathcal{P},\pi_{e}}[h(X_{t},A_{t},Y_{t})^{2}]<m$ for some constant $m$, then $\mathcal{M}_{\theta}$ will have finite bracketing numbers for all $\epsilon>0$. We place a similar assumption on the complexity of the model classes.
@@
The predictive model $f_{t}\in\sigma(\mathcal{H}_{t-1})$, and $f_{t}$ belongs to a function class $\mathcal{F}$. For this function class, there exists a single function $u_{\mathcal{F}}$ and a constant $m_{u}\in\mathbb{R}$ such that $\mathbb{E}[u_{\mathcal{F}}(X_{t},a,f_{t}(a,X_{t}))^{4}]<m_{u}$ for all $a\in\mathcal{A}$ and,
@@
for all $a\in\mathcal{A}$, $x$ in the support of $\mathcal{P}$ and $f_{t}\in\mathcal{F}$. Furthermore, we assume that there exists a constant $C_{2}$ not dependent on $t$ such thate
$\mathbb{E}\left[\left\lVert m_{\theta}(X_{t},a,f_{t}(X_{t},a))-m_{\theta^{\prime}}(X_{t},a,f_{t}(X_{t},a))\right\rVert^{2}_{2}|\mathcal{H}_{t-1}\right]\leq C_{2}\left\lVert\theta-\theta^{\prime}\right\rVert_{2}^{2}$
almost surely.
Note that a primary difference between 5 and 6 is that 6 is written in terms of expectations relative to the actual policy $\pi_{t}$ rather than $\pi_{e}$. This is necessary because 6 bounds the variance of a term related to $f_{t}$, which is dependent on history whereas 5 is bounding the complexity of a quantity that is independent of history. These assumptions allow us to prove a key lemma needed to ensure the consistency of $\hat{\theta}_{T}$ and $\tilde{\theta}_{T}$.
@@
Assume that there exists a constant $L$ not dependent on $t$ such that
$\mathbb{E}\left[\left\lVert m_{\theta}(X_{t},a,f_{t}(X_{t},a))-m_{\theta^{\prime}}(X_{t},a,f_{t}(X_{t},a))\right\rVert^{2}_{2}|\mathcal{H}_{t-1}\right]\leq L\left\lVert\theta-\theta^{\prime}\right\rVert_{2}^{2}$
almost surely. Then under Assumptions 1-4,
$\sup_{\theta\in\Theta}\frac{1}{T}\sum_{t=1}^{T}\left(R_{t}(\theta)-\mathbb{E}[R_{t}(\theta)|\mathcal{H}_{t-1}]\right)\overset{p}{\to}0.$
@@
and $\mathbb{E}_{\mathcal{P},\pi_{e}}\left[u_{m}(X_{t},A_{t},Y_{t})^{2}\right]$ is bounded.
@@
Let $\theta^{\star}$ be defined as Equation 3 and $s_{t,\theta}$ defined as in Equation 7. Assume $V_{t,\theta^{\star}}:=\mathbb{E}_{\mathcal{P},\pi_{t}}\left[s_{t,\theta^{\star}}s_{t,\theta^{\star}}^{T}|\mathcal{H}_{t-1}\right]$ is almost surely invertible and that there exists a sequence of random matrices $\{\hat{V}_{t}\}_{t=1}^{T}$ adapted to the filtration $\sigma(\mathcal{H}_{t-1})$ such that $\left\lVert\hat{V}_{t}^{-1/2}-V_{t,\theta^{\star}}^{-1/2}\right\rVert_{\text{op}}\overset{p}{\to}0$. Assume that the eigenvalues of both $\hat{V}_{t}$ and $V_{t,\theta^{\star}}$ are bounded above and below by constants $\delta_{\text{min}},\delta_{\text{max}}$. Let $\hat{\theta}_{T}$ be any estimator such that $\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\hat{\theta}_{T}}=o_{p}(1/\sqrt{T})$. Then under Assumptions 1-6,
@@
Note that letting $\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\hat{\theta}_{T}}=o_{p}(1/\sqrt{T})$ instead of exactly zero ensures that the results are still valid when using an approximate root finding algorithm instead of an exact root. As an illustrative example, we first demonstrate the application of Theorem 1 in the case of a linear model.
An illustration: correctly specified models
Let us first consider the case of correct specification, where $Y_{t}=Z_{t}^{T}\theta^{*}+\epsilon_{t}$ and $\epsilon_{t}$ are independent Gauassian noise with $\epsilon_{t}\sim N(0,\sigma_{t}^{2})$. We assume $Z_{t}\in\sigma(\mathcal{H}_{t-1})$ is dependent on user choices based on a shared history $\mathcal{H}_{t-1}:=\{Z_{1},Y_{1},...,Z_{t-1},Y_{t-1}\}$. In our framework, we can think of $Z_{t}=\phi(A_{t})$ as some deterministic function of the observed actions (e.g., one-hot encoding).
@@
Letting $\Sigma_{T}:=\sum_{t=1}^{T}Z_{t}Z_{t}^{T}$, $\lambda_{\text{min}}(\Sigma_{T})$ represent the minimum eigenvalue of $\Sigma_{T}$, and $\lambda_{\text{max}}(\Sigma_{T})$ represent the maximum eigenvalue of $\Sigma_{T}$, Lai and Wei (1982) show that a sufficient condition for the OLS estimator $\tilde{\theta}_{T}:=\Sigma_{T}^{-1}\sum_{t=1}^{T}Z_{t}Y_{t}$ to converge to $\theta$ almost surely is that $\lambda_{\text{min}}(\Sigma_{T})\overset{\text{a.s.}}{\to}\infty$ and $\frac{\log\lambda_{\text{max}}(\Sigma_{T})}{\lambda_{\text{min}}(\Sigma_{T})}\overset{\text{a.s.}}{\to}0$. However, the requirement for asymptotic normality is much stronger. They require that a deterministic series of matrices $\{B_{T}\}_{T=1}^{\infty}$ exist such that $B_{T}^{-1}(\sum_{t=1}^{T}Z_{t}Z_{t}^{T})^{1/2}\overset{p}{\to}I_{p}$.
@@
Suppose that we apply Theorem 1 in the case where $\tilde{\theta}_{T}$ is consistent but the asymptotic normality condition does not apply. We can still form a confidence interval for $\theta^{*}$ as follows. For simplicity, we assume that $f_{t}=c$ for some constant $c$ so the score function corresponds to the ordinary least squares loss without augmented components. In this case, we have $s_{t,\theta^{\star}}=-2w_{t}(Y_{t}-Z_{t}^{T}\theta^{*})Z_{t}$ with $w_{t}=\frac{\pi_{e}(A_{t})}{\mathbb{P}(A_{t}|\mathcal{H}_{t-1})}$. Moreover,
@@
$V_{t,\theta^{\star}}=\mathbb{E}\left[s_{t,\theta^{\star}}s_{t,\theta^{\star}}^{T}|\mathcal{H}_{t-1}\right]=\mathrm{Var}\left(Z_{t}^{T}(Y_{t}-\theta^{*}Z_{t}^{T})|\mathcal{H}_{t-1}\right)=\sigma_{t}^{2}\mathbb{E}\left[w_{t}^{2}Z_{t}Z_{t}^{T}|\mathcal{H}_{t-1}\right],$
@@
noting that $\text{Var}\left(Y_{t}|Z_{t}\right)=\text{Var}\left(\epsilon_{t}|\mathcal{H}_{t-1}\right)=\sigma_{t}^{2}$ and $\dot{s}_{t,\theta}=w_{t}Z_{t}Z_{t}^{T}$.
Note that we can obtain a closed form solution by solving for $\hat{\theta}_{T}$ in the equation
$0=\frac{1}{T}\sum_{t=1}^{T}V_{t,\theta^{\star}}^{-1/2}s_{t,\hat{\theta}}=\frac{1}{T}\sum_{t=1}^{T}-\left(w_{t}^{2}\sigma_{t}^{2}\mathbb{E}\left[Z_{t}Z_{t}^{T}|\mathcal{H}_{t-1}\right]\right)^{-1/2}w_{t}(Y_{t}-Z_{t}^{T}\theta^{*})Z_{t}.$
Rearranging terms yields
$\hat{\theta}_{T}=\left(\sum_{t=1}^{T}\sigma_{t}^{-1}\mathbb{E}\left[w_{t}^{2}Z_{t}Z_{t}^{T}|\mathcal{H}_{t-1}\right]^{-1/2}Z_{t}Z_{t}^{T}\right)^{-1}\left(\sum_{t=1}^{T}\sigma_{t}^{-1}w_{t}^{-1}\mathbb{E}\left[w_{t}^{2}Z_{t}Z_{t}^{T}|\mathcal{H}_{t-1}\right]^{-1/2}Z_{t}Y_{t}\right).$
The formation of confidence intervals using Theorem 1 yields,
$\frac{1}{\sqrt{T}}\sum_{t=1}^{T}w_{t}\sigma_{t}^{-1}\mathbb{E}\left[w_{t}^{2}Z_{t}Z_{t}^{T}|\mathcal{H}_{t-1}\right]^{-1/2}Z_{t}Z_{t}^{T}\left(\hat{\theta}_{T}-\theta^{\star}\right)\overset{d}{\to}N(0,I_{p}).$
In the case where the model is specified correctly, $\hat{\theta}_{T}$ is a consistent estimate of $\theta^{\star}$ for any choices of $w_{t}$ that follow 4. Therefore, the choice of $w_{t}$ becomes strictly a question of asymptotic efficiency. We can illustrate this with a few special cases
  • Suppose $\mathbb{E}\left[Z_{t}Z_{t}^{T}|\mathcal{H}_{t-1}\right]$ is independent of $\mathcal{H}_{t-1}$, then $\frac{1}{T}\sum_{t=1}^{T}Z_{t}Z_{t}^{T}$ becomes a consistent estimate for $\mathbb{E}\left[Z_{t}Z_{t}^{T}\right]$.
    • When $\sigma_{t}$ is constant across $t$, it is optimal to choose $w_{t}=1$. This reduces to the ordinary least squares solution (i.e., $\tilde{\theta}_{T}=\hat{\theta}_{T}$) and the confidence intervals become standard.
    • When $\sigma_{t}$ is not constant, it is optimal to choose $w_{t}\propto\frac{1}{\sigma_{t}^{2}}$. This reduces to the case of weighted least squares, and the confidence intervals again become standard.
  • When each component of $Z_{t}$ corresponds to an indicator variable (i.e. $Z_{t,j}=1_{A_{t}=j}$), then $\mathbb{E}\left[w_{t}^{2}Z_{t}Z_{t}^{T}|\mathcal{H}_{t-1}\right]=\text{diag}\left(w_{t}^{2}\mathbb{P}\left(Z_{t,j}=1|\mathcal{H}_{t-1}\right)\right)$. In this case, the CLT simplifies to:
    $\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\sigma_{t}^{-1/2}\text{diag}\left(\{\mathbb{P}(A_{t}=a|\mathcal{H}_{t=1})^{-1/2}\}_{a\in\mathcal{A}}\right)\left(\hat{\theta}_{T}-\theta^{\star}\right)\overset{d}{\to}N(0,I_{p}).$
Of course, this example is unrealistic because it assumes oracle knowledge of $V_{t,\theta^{\star}}=\mathbb{E}[w_{t}^{2}Z_{t}Z_{t}^{T}]$. In practice, the key challenge in applying Theorem 1 is to find a reliable method to estimate this quantity. In the next section, we tackle the problem of finding practical methods to construct a sequence of estimators $\{\hat{V}_{t}\}_{t=1}^{T}$ that can be used as plug-ins.